Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Wavefront sensing)
rocchi - 11:02 Thursday 07 January 2021 (50376) Print this report
HWS camera initialization

After the problem reported in 50321, which affected also the machine controlling the HWSs, yesterday evening I tried to re-initialize some HWS cameras.

I was successful with DET and WE CCDs (INJ had already been done by Ilaria Nardecchia), while NE resulted to be unreacheable (likely connected to the issue reported in 50368).

I profited for taking some images to check the alignment of the beams.

Figure 1 shows the image on HWS-INJ and seems OK.

Figure 2 is the beam as seen from HWS-DET. It appears to be either clipped or not well aligned, I cannot say if this is related to the activity on the detection benches (50369) or CITF optics not being well aligned. Further check is needed.

Figure 3 is the image on HWS-HR_WE, which shows only sensor noise as either the SLED is off or the beam is not reaching the CCD.

Images attached to this report
Comments to this report:
tacca - 17:49 Thursday 07 January 2021 (50382) Print this report
Following Ilaria's request, I went to check at theNE building. Both the HWS and the RH suppliers were unplugged. Since no multiplugs are present yet on the new TCS/CAL rack, I replugged the two devices on the next rack and I switched them on. Now both the RH and the HWS suppliers are ON.
rocchi - 19:57 Thursday 07 January 2021 (50384) Print this report

Thanks Matteo, the RH process has been re-started.

ballardin, rocchi - 20:48 Thursday 07 January 2021 (50383) Print this report

The HWS-HR_NE CCD has been re-initialized and is seeing a beam back (See Figure).

Images attached to this comment
Search Help
×

Warning

×